OpenStudy (anonymous):

Calculus The rate of change of the price of a vintage 1963 Chevy Corvette Coupe, in dollars per month, is modeled by the function P(t)=500cos(pi/t) for )

OpenStudy (anonymous):

\[P(t)=500\cos \left(\begin{matrix}\pi \\ t\end{matrix}\right)\] for \[0\le t \le 12\]

OpenStudy (anonymous):

The price is increasing at \(t=3\) if \(P'(3)>0\) and decreasing if \(P'(3)<0\).

OpenStudy (anonymous):

can you show me how to find the derivative to the equation/

OpenStudy (anonymous):

\[P(t)=500\cos\frac{\pi}{t}\] Apply the chain rule. The "outside" function is \(\cos\), and the "inside" function is \(\dfrac{\pi}{t}\): \[P'(t)=500\left[\frac{d}{dt}\cos\right]\frac{\pi}{t}\cdot\frac{d}{dt}\left[\frac{\pi}{t}\right]\] You won't find this notation in a textbook, so I should explain. Normally, you would write \(f(t)=500\cos t\) and \(g(t)=\dfrac{\pi}{t}\). The given function is then a composition, \(f(g(t))\). So, by the chain rule, the derivative is \[f'(g(t))\cdot g'(t)=500\left(-\sin\frac{\pi}{t}\right)\cdot \left(-\frac{\pi}{t^2}\right)\]

OpenStudy (anonymous):

So when I plug in t=3 would I use sin to find \[(-\frac{ \pi}{ t^2})\]

OpenStudy (mathmale):

No. The expression 500(−sin[π/t])⋅(−π/[t^2)]) has three separate factors, including the -500 and the sine term. Evaluate the third term separately, by letting t = 3.

OpenStudy (mathmale):

Then multiply the three (all numeric) factors together.